OpenStudy (anonymous):

how would you find the x and y intercepts of y=x^2+x-2?

OpenStudy (johnweldon1993):

Well the easiest way would to be by graphing the function, but if we cant do that we can always algebraically figure it out

OpenStudy (anonymous):

how would you solve graphically? I know the first step is to make either x or y equal 0

OpenStudy (anonymous):

** I mean algebraically

OpenStudy (johnweldon1993):

Well you are correct that is how we do it To solve for the 'y intercept' we make x = 0 \(\large y = 0^2 + 0 - 2 \) meaning that \(\large y = ?\)

OpenStudy (anonymous):

-2

OpenStudy (johnweldon1993):

Right And now the x-intercept would be making y = 0 So \[\large 0 = x^2 + x - 2\] We can factor this or just brute force through it

OpenStudy (usukidoll):

to find the y-intercept we let x = 0 to find the x-intercept we let y = 0 so when x = 0, y = -2 so that would mean that our y-intercept is at (0,-2)

OpenStudy (usukidoll):

we can factor \[\large 0 = x^2 + x - 2 \]. and then solve for x. We would have more than one solution for our x-intercepts

OpenStudy (anonymous):

thank you both!

OpenStudy (usukidoll):

so we need to focus on the last term 2 and the second term 1 there's only one combination 2 x 1 = 2 (so we got the last term) we need to use subtraction to obtain one 2-1 =1 so we have \[\large 0 = (x+2)(x-1)\]

OpenStudy (usukidoll):

now we just split this up and solve for x in each case \[\large 0 = x-1\],\[\large 0 = x+2 \]

OpenStudy (anonymous):

so xint= -2 and 1 ?

OpenStudy (usukidoll):

yes our x = -2 and x = 1 so our x-intercepts are (-2,0) and (1,0) because we need y = 0 to obtain the x-intercepts.

OpenStudy (anonymous):

first put y=0 so your equation will be0=x^2+x−2 then you will make it (x-1)(x+2)=0 so it will lead that x is equal to 1 and -2